OpenStudy (anonymous):

Evaluate the expression – 4a – 5b – 9 for a = 3 and b = –2 in a test need help now

OpenStudy (helder_edwin):

since a=3 and b=-2 then \[ \large -4a-5b-9=-4(3)-5(-2)-9= \]

OpenStudy (anonymous):

A.-31 b.25 c.-11 d.13

OpenStudy (helder_edwin):

do the arithmetic!

OpenStudy (anonymous):

i tried and i want your opinion

OpenStudy (anonymous):

There is no opinion. You either did it the right way or the wrong way. Show your work and tell us what you got

OpenStudy (helder_edwin):

what did u get?

OpenStudy (anonymous):

-31

OpenStudy (anonymous):

Can you show us your work?

OpenStudy (anonymous):

what did u get

OpenStudy (anonymous):

I haven't done it yet.

OpenStudy (anonymous):

truy

OpenStudy (anonymous):

I want to see your work first before I try.

OpenStudy (helder_edwin):

\[ \large =-12+10-9 \]

OpenStudy (helder_edwin):

u can do it! just focus

OpenStudy (anonymous):

-11

OpenStudy (helder_edwin):

yes. did u get it?

OpenStudy (anonymous):

yes mathmatical

OpenStudy (helder_edwin):

do u understand why -12 and why +10?

OpenStudy (anonymous):

yeah

OpenStudy (helder_edwin):

ok

OpenStudy (anonymous):

by

OpenStudy (helder_edwin):

bye

OpenStudy (anonymous):

Oh snap this guy is taking a test @helder_edwin We are not suppose to help him.

OpenStudy (anonymous):

i like penises